A jet flew at an average speed of 640 mph from Los Angeles to Chicago. Because of head winds the jet averaged only 512 mph on the return trip, and the return trip took 48 minutes longer. How many hours was the flight from Chicage to Los Angeles? How far is it from Chicago to Los Angeles?
Answer by Paul(988)

48 minutes longer mean add to the support of the second number 512, and would mean do the opposite of the first number.
640(t-48)=512(t+48)
640t-30720=512t+24576
128t=55296
t=432
Hence, Chicago is 432 miles apart from Los Angeles.